# One Step is Enough: Multi-Agent Reinforcement Learning based on One-Step Policy Optimization for Order Dispatch on Ride-Sharing Platforms

### v1

@@ Abstract
On-demand ride-sharing platforms face the fundamental challenge of dynamically bundling passengers with diverse origins and destinations and matching them with vehicles in real time, all under significant uncertainty. Recently, Multi-Agent Reinforcement Learning (MARL) has emerged as a promising solution for this problem, leveraging decentralized learning to address the curse of dimensionality caused by the large number of agents in the ride-hailing market and the resulting expansive state and action spaces. However, conventional MARL-based ride-sharing approaches heavily rely on the accurate estimation of Q-values or V-values, which becomes problematic in large-scale, highly uncertain environments. Specifically, most of these approaches adopt an independent paradigm, exacerbating this issue, as each agent treats others as part of the environment, leading to unstable training and substantial estimation bias in value functions. To address these challenges, we propose two novel alternative methods that bypass value function estimation. First, we adapt Group Relative Policy Optimization (GRPO) to ride-sharing, replacing the Proximal Policy Optimization (PPO) baseline with the group average reward to eliminate critic estimation errors and reduce training bias. Second, inspired by GRPO’s full utilization of group reward information, we customize the PPO framework for ride-sharing platforms and show that, under a homogeneous fleet, the optimal policy can be trained using only one-step rewards—a method we term One-Step Policy Optimization (OSPO). Experiments on a real-world Manhattan ride-hailing dataset demonstrate that both GRPO and OSPO achieve superior performance across most scenarios, efficiently optimizing pickup times and the number of served orders using simple Multilayer Perceptron (MLP) networks. Furthermore, OSPO outperforms GRPO in all scenarios, attributed to its elimination of bias caused by the bounded simulation time of GRPO. Our code, trained models, and processed data are publicly available at the anonymous repository: https://github.com/RS2002/OSPO.
Introduction
@@
The widespread adoption of on-demand ride-sharing platforms has fundamentally transformed urban transportation systems, offering scalable and sustainable mobility solutions for modern cities. By enabling multiple passengers with similar itineraries to share a single vehicle, ride-sharing platforms alleviate congestion, reduce vehicle kilometers traveled, and decrease urban fleet sizes, thereby addressing pressing challenges such as traffic, emissions, and inefficient resource utilization (Wang and Yang 2019; Santi et al. 2014). However, managing such platforms presents significant operational complexity. Passengers arrive randomly over time, each with unique origins and destinations, and the platform must dynamically determine not only how to bundle new orders together, but also how to assign each bundle to a suitable vehicle. These dispatching decisions must take into account the current routing of each vehicle, the destinations of onboard passengers, and the service quality requirements of all users. Critically, all decisions must be made in real time, under uncertainty regarding future demand and supply, making the order dispatch problem a central and challenging task for ride-sharing platforms.
@@
Recently, reinforcement learning (RL) methods have shown great promise in addressing the order dispatch problem on ride-sharing platforms. Although ride-sharing is fundamentally a centralized task—requiring the platform to aggregate information from all drivers and orders to make globally optimal decisions—the massive state and action spaces encountered in realistic settings (which can reach millions or billions (Sivagnanam et al. 2024)) pose a serious curse of dimensionality, rendering single-agent RL approaches intractable. As a result, most recent works adopt the Multi-Agent Reinforcement Learning (MARL) framework, where each driver is treated as an individual agent, and a global controller coordinates the agents to optimize the overall system objectives. MARL methods are typically categorized into three types: Decentralized Training with Decentralized Execution (DTDE), Centralized Training with Decentralized Execution (CTDE), and Centralized Training with Centralized Execution (CTCE) (Jin et al. 2025). Most existing ride-sharing approaches follow the DTDE paradigm, since centralized methods still face scalability limitations similar to single-agent RL, and are thus difficult to apply in large-scale scenarios with hundreds or thousands of agents. However, DTDE methods rely on the independent assumption that each agent acts based only on its local information and treats other agents as part of the environment. This often leads to unstable training and poor cooperation among agents, as well as large estimation errors in Q-values and V-values (Hu, Feng, and Li 2025), ultimately resulting in suboptimal and inefficient learning.
@@
To address this problem, we propose two multi-agent policy optimization methods that do not require estimating Q-values or V-values. First, we adapt Group Relative Policy Optimization (GRPO) (Shao et al. 2024) for the ride-sharing task, replacing the V-value baseline in Proximal Policy Optimization (PPO) (Schulman et al. 2017) with the group average reward. This modification significantly reduces the computational resources required for training and minimizes bias by eliminating neural network estimation errors associated with V-values and Q-values. Furthermore, inspired by the baseline replacement strategy in GRPO, we introduce a more efficient method called One-Step Policy Optimization (OSPO), which can be trained using only one-step rewards by leveraging the homogeneous characteristics of agents in the ride-sharing task. Both methods are validated through case studies using real-world ride-hailing data and demonstrate superior performance compared to state-of-the-art approaches. The main contributions of this paper are summarized as follows:
@@
  • We introduce a novel ride-sharing framework based on GRPO. By replacing the baseline in PPO with group average reward, our method eliminates the need to estimate V-values. It successfully improves training performance and stability by removing the influence of estimation bias. To the best of our knowledge, this is the first successful application of GRPO in a non-LLM and MARL context, specifically in ride-sharing platforms.
@@
  • Inspired by GRPO, we further propose a more efficient policy optimization method named OSPO, where we derive that, according to the agent homogeneity property in ride sharing task, the advantage function can be replaced by the group relative reward within only one step. To the best of our knowledge, it is the simplest and most efficient RL method for the ride-sharing task. Furthermore, we also specify the conditions under which OSPO can be applied in identical MARL scenarios.
@@
Figure 1: The training process consists of four steps: (1) Calculate the matching probability for each driver-order pair; (2) Assign orders by maximizing the probability score; (3) Calculate the group advantage; (4) Train the network using PPO-style policy gradients.
Preliminary
Group Relative Policy Optimization (GRPO)
@@
With the success and popularity of Reinforcement Learning from Human Feedback (RLHF), RL-based LLM post-training has garnered increasing attention. This has also led to the emergence of many new RL methods, such as Direct Preference Optimization (DPO) (Rafailov et al. 2023), Reinforce Leave-One-Out (RLOO) (Ahmadian et al. 2024), and GRPO (Shao et al. 2024). Among these, GRPO is noted for its simple format, high efficiency, and promising performance, which are core factors contributing to the success of DeepSeekMath (Shao et al. 2024) and DeepSeek-V3 (Liu et al. 2024).
@@
Serving as a PG (Sutton et al. 1999) based method, the optimization objective of GRPO has a similar format to PPO(Schulman et al. 2017), expressed as:
$\text{J}_{GRPO}(\theta)=\mathbb{E}_{s,u\sim\pi_{\theta^{-}}}\left[\min\left\{\frac{\pi_{\theta}(s)}{\pi_{\theta^{-}}(s)}\hat{\text{A}}_{\pi_{\theta}}(s,u),\text{CLIP}\right.\right.$ (1)
@@
$\left.\left(\frac{\pi_{\theta}(s)}{\pi_{\theta^{-}}(s)},1-\epsilon,1+\epsilon^{\prime}\right)\hat{\text{A}}_{\pi_{\theta}}(s,u)\right\}\Bigg{]}-\beta\text{KL}(\pi_{\theta}\|\pi_{ref})\ ,$
@@
where $\pi$ denotes the policy, $\theta$ and $\theta^{-}$ represent the current network parameters and the parameters used for experience collection, respectively, $s$ and $u$ denote the state and action, $\pi_{ref}$ is the reference policy (i.e., the initial policy before post-training), serving as a normalization term to prevent the policy from diverging too much from the reference, $\hat{\text{A}}_{\pi_{\theta}}$ represents the advantage function of policy $\pi_{\theta}$, and $\beta,\epsilon,\epsilon^{\prime}$ are hyper-parameters We follow the improvement in (Yu et al. 2025), which suggests setting a higher value for $\epsilon^{\prime}$ than for $\epsilon$ to encourage exploration.
@@
Problem Setup for Ride-sharing Order Dispatch
In this paper, we consider the optimal order assignment task of an on-demand ride-sharing platform with $n$ homogeneous drivers (agents), each equipped with a vehicle of capacity $c$. At each time step $t$, there will be $m_{t}$ newly arrived orders from customers in different locations. The platform must dynamically determine not only how to bundle new orders together, but also how to assign each bundle to a suit- able vehicle. These dispatching decisions must take into account the current routing of each vehicle, the destinations of onboard passengers, and the service quality requirements of all users. Specifically, we model this process as a Multi-Agent Markov Decision Process (MAMDP) (Littman 1994), defined as $<n,S,U,P,R,\gamma>$, where $n$ represents the number of agents, and $\gamma$ represents the discount factor. The global state $S$ can be written as $S=(s^{o},s_{1},s_{2},\ldots,s_{n})$, where $s^{o}$ represents the global state (i.e. information of orders to be assigned), visible to all agents, and $s_{i}$ represents the state of agent $i$. In our ride-sharing task, we set $s^{o}$ to consist of the OD and arrival times of all unconfirmed orders, while $s_{i}$ consists of the current location, remaining capacity, the cumulative reward since the episode started, the average cumulative reward of all agents, and the information of all onboard orders, including their destination, estimated total delivery time, and remaining delivery time. The global action $U$ is the joint action of all agents, expressed as $U=(u_{1},u_{2},\ldots,u_{n})$. At time $t$, the action $u_{i}$ is a $w_{t}$-dimensional 0-1 vector, where $u_{i,j}=1$ indicates that order $j$ is assigned to agent $i$, and $w_{t}$ is the total number of orders to be assigned at time $t$. The transition probability function $P$ is not modeled explicitly in model-free RL methods. For the reward function $\text{r}(s_{i},u_{i})$, we leave the detailed design in appendix. Since ride-sharing is a fully cooperative task, the global reward can be expressed as the sum of rewards from each individual agent, written as $\text{R}(S,U)=\sum_{i=1}^{n}\text{r}(s_{i},u_{i})$.
Methodology
In this section, we first describe how we adapt GRPO for the multi-agent ride-sharing scenario, and analyze the theoretical foundations that support its applicability. We then introduce our proposed OSPO method and specify the conditions under which it can be effectively applied.
GRPO for Ride Sharing
The workflow of GRPO for ride sharing is shown in Fig. 1, following a design similar to previous SAC-based work (Enders et al. 2023). Firstly, we require the policy network $\pi_{\theta}$ to directly output the action probability of each agent, expressed as:
@@
Considering the action constraint that an order cannot be assigned to multiple agents, we cannot simply choose the action independently for each agent. Instead, we view the action probabilities as scores and use bipartite matching to maximize the assignment score, following (Enders et al. 2023; Hu, Feng, and Li 2025):
@@
where $u_{i,j,t}$ indicates whether agent $i$ is assigned order $j$ at time $t$ (with 1 indicating assignment and 0 indicating no assignment), and $y_{i,j,t}$ denotes the score $p_{i,j,t}$ of agent $i$ choosing order $j$ at time $t$ (with $y_{i,j,t}=-\infty$ for all unavailable workers at time $t$). The set =$\mathcal{I}=\{1,2,\ldots,n\}$ represents the agents, and the set $\mathcal{J}_{t}=\{1,2,\ldots,w_{t}\}$ represents the orders to be assigned. Constraint (5b) ensures that an order can be assigned to at most one agent, while constraint (5c) guarantees that each agent is assigned at most one order. After that, we can obtain the reward according to Eq. 13 and then optimize the policy by maximizing Eq. 1.
To make GRPO work in our homogeneous cooperative MARL scenario, only two minor changes should be made to the original objective function in the SARL scenario. First, when calculating the advantage function, we use the rewards of all agents over all steps to compute the mean value $\mu$ and standard deviation $\sigma$ in Eq. 3. (For stability, we propose to ignore samples with a reward of 0 when calculating $\mu$ and $\sigma$. This is feasible since those no-assignment samples do not correspond to a real action and therefore cannot be used for training.) Second, for the KL normalization term in Eq. 1, since we do not have a reference policy like the LLM post-training, we propose to regularly track policy performance and retain the best checkpoint. This best policy serves as the reference policy, helping to ensure that the current policy does not deviate significantly from the best one. A detailed algorithm is provided in appendix.
Next, we analyze why our modification to GRPO is valid, allowing the transfer from SARL to MARL. For the advantage function (Eq. 3) in GRPO, we can express it as:
$\hat{\text{A}}_{\pi}(s_{t},u_{t})=\frac{1}{\sigma}\left(\sum_{\tau\in K_{t}}\gamma^{\tau}\text{R}(s_{t+\tau},u_{t+\tau})-\sum_{\tau\in K_{t}}\gamma^{\tau}\mu\right)\ .$ (6)
For the term $\sum_{\tau\in K_{t}}\gamma^{\tau}\text{R}(s_{t+\tau},u_{t+\tau})$, we can view it as an estimation of $\text{Q}_{\pi}(s_{t},u_{t})$, similar to what we do in PPO (Schulman et al. 2017). For the term $\sum_{\tau\in K_{t}}\gamma^{\tau}\mu$, we can write it as $\sum_{\tau\in K_{t}}\gamma^{\tau}(\mathbb{E}_{s,u\sim\pi}[\text{R}(s,u)]+\xi)$, where $\xi$ is the estimation error, which will reduce to 00 as the number of sampled trajectories increases (i.e., as the group becomes larger).
The original application scenario of GRPO is in LLM post-training, where the reward function is designed through a process supervision approach (Wang et al. 2023). An important feature of such a reward function is that each step reward has a similar range, leading to: $\mathbb{E}_{s_{t},u_{t}\sim\pi}[\text{R}(s_{t},u_{t})]=\mathbb{E}_{s,u\sim\pi}[\text{R}(s,u)]+\xi_{t}$, where $\xi_{t}$ is a small residual and $\mathbb{E}_{t}[\xi_{t}]=0$. Thus, we have:
$\sum_{\tau\in K_{t}}\gamma^{\tau}(\mathbb{E}_{s,u\sim\pi}[\text{R}(s,u)]+\xi)$ (7)

### v2

@@ Abstract
Order dispatch is a critical task in ride-sharing systems with Autonomous Vehicles (AVs), directly influencing efficiency and profits. Recently, Multi-Agent Reinforcement Learning (MARL) has emerged as a promising solution to this problem by decomposing the large state and action spaces among individual agents, effectively addressing the Curse of Dimensionality (CoD) in transportation market, which is caused by the substantial number of vehicles, passengers, and orders. However, conventional MARL-based approaches heavily rely on accurate estimation of the value function, which becomes problematic in large-scale, highly uncertain environments. To address this issue, we propose two novel methods that bypass value function estimation, leveraging the homogeneous property of AV fleets. First, we draw an analogy between AV fleets and groups in Group Relative Policy Optimization (GRPO), adapting it to the order dispatch task. By replacing the Proximal Policy Optimization (PPO) baseline with the group average reward-to-go, GRPO eliminates critic estimation errors and reduces training bias. Inspired by this baseline replacement, we further propose One-Step Policy Optimization (OSPO), demonstrating that the optimal policy can be trained using only one-step group rewards under a homogeneous fleet. Experiments on a real-world ride-hailing dataset show that both GRPO and OSPO achieve promising performance across all scenarios, efficiently optimizing pickup times and the number of served orders using simple Multilayer Perceptron (MLP) networks. Furthermore, OSPO outperforms GRPO in all scenarios, attributed to its elimination of bias caused by the bounded time horizon of GRPO. Our code, trained models, and processed data are provided at https://github.com/RS2002/OSPO.
Machine Learning, ICML
1 Introduction
The widespread adoption of on-demand ride-sharing platforms has fundamentally transformed urban transportation systems, offering scalable and sustainable mobility solutions for modern cities. By enabling multiple passengers with similar itineraries to share a single vehicle, ride-sharing platforms alleviate congestion, reduce vehicle kilometers traveled, and decrease urban fleet sizes, thereby addressing pressing challenges such as traffic, emissions, and inefficient resource utilization (wang2019ridesourcing; santi2014quantifying). Recently, with the development of Autonomous Vehicle (AV) technology, systems like Robotaxi have been introduced by various companies, including Baidu, Didi, and Zoox. These systems not only offer lower operating costs through the AV fleet but also enable better management, as the fleet is fully controlled rather than relying on human drivers.
@@ 1 Introduction
However, managing such platforms presents significant operational complexity. Passengers arrive randomly over time, each with unique origins and destinations, and the platform must dynamically determine not only how to bundle new orders together, but also how to assign each bundle to a suitable vehicle. These dispatching decisions must take into account the current routing of each vehicle, the destinations of onboard passengers, and the service quality requirements of all users. Critically, all decisions must be made in real time, under uncertainty regarding future demand and supply, making the order dispatch problem a central and challenging task for ride-sharing platforms with AVs.
@@
Recently, Reinforcement Learning (RL) methods have shown great promise in addressing the order dispatch problem on ride-sharing platforms. Although order dispatch is fundamentally a centralized task—requiring the platform to aggregate information from all vehicles and orders to make globally optimal decisions—the massive state and action spaces encountered in realistic settings (which can reach millions or billions (sivagnanam2024multi)) pose a serious Curse of Dimensionality (CoD), rendering Single-Agent Reinforcement Learning (SARL) approaches intractable. As a result, most recent works adopt the Multi-Agent Reinforcement Learning (MARL) framework, where each vehicle is treated as an individual agent, and a global controller coordinates the agents to optimize the overall system objectives. MARL methods are typically categorized into three types: Decentralized Training with Decentralized Execution (DTDE), Centralized Training with Decentralized Execution (CTDE), and Centralized Training with Centralized Execution (CTCE) (jin2025comprehensive). Most existing ride-sharing approaches follow the DTDE paradigm, since centralized methods still face scalability limitations similar to SARL, and are thus difficult to apply in large-scale scenarios with hundreds or thousands of agents. However, DTDE methods rely on the independent assumption that each agent acts based only on its local observation and treats other agents as part of the environment. This often leads to unstable training and poor cooperation among agents, as well as large estimation errors in Q-values and V-values (hu2025bmg), ultimately resulting in suboptimal and inefficient learning.
@@
To address this problem, we propose two novel multi-agent policy optimization methods that leverage the homogeneous characteristics of agents in the ride-sharing system with AVs, eliminating the need to estimate Q-values or V-values. First, we adapt Group Relative Policy Optimization (GRPO) (GRPO) for the order dispatch task, replacing the V-value baseline in Proximal Policy Optimization (PPO) (PPO) with the group average reward-to-go. This modification significantly reduces the computational resources required for training and minimizes bias by eliminating neural network estimation errors associated with V-values and Q-values. Furthermore, inspired by the baseline replacement strategy in GRPO, we introduce a more efficient method called One-Step Policy Optimization (OSPO), which can be trained using only one-step group rewards. Both methods are validated through case studies using real-world ride-hailing data and demonstrate promising performance with high efficiency. The main contributions of this paper are summarized as follows:
@@
  • We introduce a novel order dispatch method for ride-sharing system with AVs based on GRPO. By replacing the baseline in PPO with group average reward-to-go, our method eliminates the need to estimate V-values. It successfully improves training performance and stability by removing the influence of estimation bias. To the best of our knowledge, this is the first successful application of GRPO in a non-LLM and MARL context, specifically in ride-sharing systems.
@@
  • Inspired by GRPO, we further propose a more efficient policy optimization method named OSPO, where we derive that, according to the agent homogeneity property in AV-based ride-sharing system, the advantage function can be replaced by the group relative reward within only one step. To the best of our knowledge, it is the simplest and most efficient RL method for the order dispatch task. Furthermore, we also specify the conditions under which OSPO can be applied in identical MARL scenarios.
@@
2 Preliminary
2.1 Problem Setup
In this paper, we consider the optimal order assignment task of an ride-sharing platform with $n$ homogeneous AVs (agents), each equipped with capacity $c$. At each time step $t$, there will be $m_{t}$ newly arrived orders from passengers in different locations. The platform must dynamically determine how to assign each order to a suitable vehicle in real-time. The dispatching decisions must take into account the current routing of each vehicle, the destinations of onboard passengers, and the service quality requirements of all users. Specifically, we model this process as a Multi-Agent Markov Decision Process (MAMDP) (littman1994markov), defined as $n,S,U,\mathcal{T},R,\gamma>$, where $n$ represents the number of agents, and $\gamma$ represents the discount factor. The global state $S$ can be written as $S=(s^{o},s_{1},s_{2},\ldots,s_{n})$, where $s^{o}$ represents the global state (i.e. information of orders to be assigned), visible to all agents, and $s_{i}$ represents the state of agent $i$. In our order dispatch task, we set $s^{o}$ to consist of the OD and arrival times of all unconfirmed orders, while $s_{i}$ consists of the current location, remaining capacity, the cumulative reward since the episode started, the average cumulative reward of all agents, and the information of all onboard orders, including their destination, estimated total delivery time, and estimated remaining delivery time. The global action $U$ is the joint action of all agents, expressed as $U=(u_{1},u_{2},\ldots,u_{n})$. At time $t$, the action $u_{i}$ is a $w_{t}$-dimensional 0-1 vector, where $u_{i,j}=1$ indicates that order $j$ is assigned to agent $i$, and $w_{t}$ is the total number of orders to be assigned at time $t$. The transition probability function $\mathcal{T}$ is not modeled explicitly in model-free RL methods. At last, the reward function is shaped following (hu2025bmg):
[…]
where $\beta_{1}$ to $\beta_{5}$ are non-negative weights representing the platform’s valuation of each term, $p^{in}_{i}$ and $p^{out}_{i}$ represent the income from passengers and the vehicle operational costs, respectively. The variables $\chi_{i}$ and $\rho_{i}$ represent the number of en-route orders that will exceed their scheduled time and the additional travel time of all en-route orders when the assigned order is added to the scheduled route of agent $i$, respectively. This reward function is designed to comprehensively consider the interests of the platform and passengers, mimicking the operation of a real-world ride-sharing platform. It is important to emphasize that $p^{in}_{i}$ and $p^{out}_{i}$ are calculated based on the order distance and the additional travel distance for the agent, respectively. When calculating travel time, we will utilize the Traveling Salesman Problem (TSP) to optimize the agent’s route. Since order dispatch is a fully cooperative task, the global reward can be expressed as the sum of rewards from each individual agent, written as $\text{R}(S,U)=\sum_{i=1}^{n}\text{r}(s_{i},u_{i})$.
2.2 Homogeneous Property in Ride-Sharing System
In this paper, we leverage the homogeneous property of agents (AVs), which is foundational to our proposed methods. This homogeneity can be understood from both the task and the MARL perspectives. From the task perspective, within the order dispatch system, the primary differences between agents are their spatial locations and the assigned en-route orders. The platform itself is agnostic to individual agent identity. From the MARL perspective, all agents share an identical structure for their observation space, action space, and reward function. Furthermore, the state transition dynamics are governed by the same environmental rules. This makes the order dispatch task a fully cooperative one for the platform, where agents can share a common policy. This homogeneous property is also a key enabler for parameter and experience sharing, a common practice in prior works (al2019deeppool; enders2023hybrid; hu2025bmg).
Considering a multi-agent system, the state value function among agents exhibits the following relationship:
$V_{\pi}(s_{1,t})+\epsilon_{1,t}=V_{\pi}(s_{2,t})+\epsilon_{2,t}=\cdots$ (2)
$=V_{\pi}(s_{n,t})+\epsilon_{n,t}=\mathbb{E}_{i\in\mathcal{I},s_{i,t}\sim\pi}[V_{\pi}(s_{i,t})]\ ,$
where $\epsilon_{i,t}$ is a residual term, defined as $\epsilon_{i,t}:=\mathbb{E}_{i\in\mathcal{I},s_{i,t}\sim\pi}[V_{\pi}(s_{i,t})]-V_{\pi}(s_{i,t})$, and $\mathcal{I}=\{1,2,3,\ldots,n\}$. Given the homogeneous property among agents in our scenario, the residual term $\epsilon_{i,t}$ is very small. This aligns with the principles of Mean Field MARL (yang2018mean), which has also been adapted for order dispatch tasks in ride-sharing systems (li2019efficient).
2.3 Group Relative Policy Optimization (GRPO)
@@
With the success and popularity of Reinforcement Learning from Human Feedback (RLHF), RL-based LLM post-training has garnered increasing attention. This has also led to the emergence of many new RL methods, such as Direct Preference Optimization (DPO) (rafailov2023direct), Reinforce Leave-One-Out (RLOO) (ahmadian2024back), and GRPO (GRPO). Among these, GRPO is noted for its simple format, high efficiency, and promising performance, which are core factors contributing to the success of DeepSeekMath (GRPO) and DeepSeek-V3 (liu2024deepseek).
@@
Serving as a PG (sutton1999policy) based method, the optimization objective of GRPO has a similar format to PPO (PPO), expressed as:
[…]
@@
$\left.\left(\frac{\pi_{\theta}(u|s)}{\pi_{\theta^{-}}(u|s)},1-\epsilon,1+\epsilon^{\prime}\right)\hat{\text{A}}_{\pi_{\theta}}(s,u)\right\}\Bigg]-\beta\text{KL}(\pi_{\theta}\|\pi_{ref})\ ,$
@@
where $\pi$ denotes the policy, $\theta$ and $\theta^{-}$ represent the current network parameters and the parameters used for experience collection, respectively, $s$ and $u$ denote the state and action, $\pi_{ref}$ is the reference policy (i.e., the initial policy before post-training), serving as a normalization term to prevent the policy from diverging too much from the reference, $\hat{\text{A}}_{\pi_{\theta}}$ represents the advantage function of policy $\pi_{\theta}$, and $\beta,\epsilon,\epsilon^{\prime}$ are hyper-parameters We follow the improvement in (yu2025dapo), which suggests setting a higher value for $\epsilon^{\prime}$ than for $\epsilon$ to encourage exploration.
@@
3 Methodology
In this section, we first introduce how to adapt policy-based MARL methods for the order assignment task in the ride-sharing system (operation stage). Then, we propose our GRPO and OSPO methods as optimization solutions (training stage), accompanied by detailed theoretical analysis.
3.1 Policy-Based Method for Order Dispatch in Ride-Sharing System
In policy-based deep RL, at each time step $t$, the agent directly generates the probabilities of each action based on the observation:
@@ 3.1 Policy-Based Method for Order Dispatch in Ride-Sharing System
However, in the order dispatch task, we cannot assign the same order to multiple agents, making it infeasible to directly sample actions as $u_{i,t}\sim p_{i,t}$. To address this problem, we follow previous policy-based ride-sharing methods (enders2023hybrid; hoppe2024global) by viewing the probabilities as scores and using bipartite matching to maximize the assignment score:
@@
where $u_{i,j,t}$ indicates whether agent $i$ is assigned to order $j$ at time $t$ (with 1 indicating assignment and 0 indicating no assignment), and $y_{i,j,t}$ denotes the score $p_{i,j,t}$ of agent $i$ choosing order $j$ at time $t$ (with $y_{i,j,t}=$-$\infty$ for all unavailable workers at time $t$). The set $\mathcal{I}=\{1,2,\ldots,n\}$ represents the agents, and the set $\mathcal{J}_{t}=\{1,2,\ldots,w_{t}\}$ represents the orders to be assigned. Constraint (7b) ensures that an order can be assigned to at most one agent, while constraint (7c) guarantees that each agent is assigned at most one order. This process is illustrated in the first two steps of Fig. 1.
3.2 GRPO Solution
To motivate the adaptation of GRPO to our order dispatch task, we first analyze its theoretical underpinnings. For the advantage function (Eq. 5) in GRPO, we can express it as:
$\hat{\text{A}}_{\pi}(s_{t},u_{t})$ $=\sum_{\tau\in K_{t}}\gamma^{\tau}\frac{\text{R}(s_{t+\tau},u_{t+\tau})-\mu}{\sigma}$ (8)
$=\frac{1}{\sigma}\left(\sum_{\tau\in K_{t}}\gamma^{\tau}\text{R}(s_{t+\tau},u_{t+\tau})-\sum_{\tau\in K_{t}}\gamma^{\tau}\mu\right)\ .$
For the term $\sum_{\tau\in K_{t}}\gamma^{\tau}\text{R}(s_{t+\tau},u_{t+\tau})$, we can view it as an estimation of $\text{Q}_{\pi}(s_{t},u_{t})$, similar to what we do in PPO (PPO). For the term $\sum_{\tau\in K_{t}}\gamma^{\tau}\mu$, we can write it as $\sum_{\tau\in K_{t}}\gamma^{\tau}(\mathbb{E}_{s,u\sim\pi}[\text{R}(s,u)]+\xi)$, where $\xi$ is the estimation error, which will reduce to 0 as the number of sampled trajectories increases (i.e., as the group becomes larger).
The original application scenario of GRPO is in LLM post-training, where the reward function is designed through a process supervision approach (wang2023math). An important feature of such a reward function is that each step reward has a similar range, leading to: $\mathbb{E}_{s_{t},u_{t}\sim\pi}[\text{R}(s_{t},u_{t})]=\mathbb{E}_{s,u\sim\pi}[\text{R}(s,u)]+\xi_{t}$, where $\xi_{t}$ is a small residual and $\mathbb{E}_{t}[\xi_{t}]=0$. Thus, we have:
$\sum_{\tau\in K_{t}}\gamma^{\tau}\text{R}(s_{t+\tau},u_{t+\tau})=\sum_{\tau\in K_{t}}\gamma^{\tau}(\mathbb{E}_{s,u\sim\pi}[\text{R}(s,u)]+\xi)$ (9)